A skater of mass 76.5 kg initially moves in a straight line at a speed of 4.50 m/s. The skater approaches a child of mass 41.0 kg, whom he lifts on his shoulders. Assuming there are no external horizontal forces, what is the skater's final velocity?
Answer #1

Initial momentum = Mv = 76.5 * 4.5 = 344.25

FInal momentum = ( 76.5 + 41) * v

Equating both we have

117.5 v = 344.25

v = 2.93 m/s
Answer #2

Let the final speed of skater with child is v

Now , we will use conservation of momentum here , as there is no external force acting

76.5 * 4.50 + 41 * 0 = (71.5 + 41)* v

solving for v

v = 2.93 m/s

the final speed of skater is 2.93 m/s